OpenStudy (anonymous):

factor the polynomial complelely 99u2-704v2

OpenStudy (anonymous):

99u2-704v2 = (3sqrt(11) u +sqrt(70) v)(3sqrt(11) u -sqrt(70) v)

OpenStudy (anonymous):

\[99u^2 - 704v^2\]\[=11(9u^2 - 64v^2)\]\[=11(3u + 8v)(3u - 8v)\]